OpenStudy (anonymous):

John's daily working hours have increased from 8 hours to 10 hours but his hourly pay has decreased by 25%. Find the percentage changes in John's daily income.

OpenStudy (anonymous):

Let's call his initial hourly wage \(\bf x\). Then his initial daily salary was \(\bf 8x\). Now his hourly wage is \(\bf x-0.25x=0.75x\). Hence his new daily income is \(\bf .75x *10=7.5x\). Therefore his pay difference is \(\bf 8x-7.5x=0.5x\). Dividing the pay difference by the original daily income, we obtain the percentage by which his pay decreased in comparison to his original pay:\[\bf \frac{ 0.5x }{8x }=\frac{ 1/2 }{ 8 }=\frac{ 1 }{ 16 }\]Hence new pay has decreased by 1/16th of his original pay.

OpenStudy (anonymous):

thanks